OpenStudy (anonymous):

*********FAN AND MEDAL******* 2x + 3y = 30... Find 3 diffrent combinations of x and y's solution (I already have one which is x = 6, and y = 6. PLEASE HELP ME!!!

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@quarkine come on can you help me out Sasuke?

OpenStudy (anonymous):

@dusty any ideas on this?

OpenStudy (anonymous):

@amistre64 @ash2326 @johnweldon1993 @mathstudent55

OpenStudy (mertsj):

You want me to help you?

OpenStudy (anonymous):

yes, if you can

OpenStudy (mertsj):

Pick a number.

OpenStudy (anonymous):

3

OpenStudy (mertsj):

Multiply that number by 2

OpenStudy (anonymous):

6

OpenStudy (mertsj):

Subtract that answer from 30

OpenStudy (anonymous):

24

OpenStudy (mertsj):

Divide that answer by 3

OpenStudy (anonymous):

8

OpenStudy (mertsj):

There you go. If x = 3, y = 8

OpenStudy (mertsj):

Pick another number

OpenStudy (anonymous):

5

OpenStudy (mertsj):

Multiply that number by 2

OpenStudy (anonymous):

10

OpenStudy (mertsj):

Subtract that answer from 30

OpenStudy (anonymous):

20

OpenStudy (mertsj):

Divide that answer y 3

OpenStudy (anonymous):

6.66666

OpenStudy (mertsj):

20/3 will do just as well. There you go. If x = 5, y = 20/3

OpenStudy (mertsj):

How many more answers do you need?

OpenStudy (anonymous):

It is for a realy long word problem in math, and I need whole numbers. I just simplified it for the question on here

OpenStudy (anonymous):

and just 1

OpenStudy (amistre64):

2x + 3y = 30 3 = 2(1)+1 ; gcd(3,2) = 1 2 = 1(2)+0 0 0 1 0-1(1) = -1 ; x' = -1 1 1 0 1-0(1) = 1 ; y' = 1 1 = 2x' + 3y' 30 = 2(30x') + 3(30y') X = 30x' + 3n Y = 30y' - 2n should solve for all the integer solutions

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=2%28-30%2B3n%29%2B3%2830-2n%29 just to be sure ;)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

2x + 3y = 30 represent a line in the x-y plane.so any point on the line will satisfy this condition.So find the solution graphically .Or solve as y=(30-2x)/3 .Put any value in x and you will get a y. hence you can get as many pair as you like..